# GenStereo: Towards Open-World Generation of Stereo Images and Unsupervised Matching

### v1

@@ Abstract
Stereo images are fundamental to numerous applications, including extended reality (XR) devices, autonomous driving, and robotics. Unfortunately, acquiring high-quality stereo images remains challenging due to the precise calibration requirements of dual-camera setups and the complexity of obtaining accurate, dense disparity maps. Existing stereo image generation methods typically focus on either visual quality for viewing or geometric accuracy for matching, but not both. We introduce GenStereo, a diffusion-based approach, to bridge this gap. The method includes two primary innovations (1) conditioning the diffusion process on a disparity-aware coordinate embedding and a warped input image, allowing for more precise stereo alignment than previous methods, and (2) an adaptive fusion mechanism that intelligently combines the diffusion-generated image with a warped image, improving both realism and disparity consistency. Through extensive training on 11 diverse stereo datasets, GenStereo demonstrates strong generalization ability. GenStereo achieves state-of-the-art performance in both stereo image generation and unsupervised stereo matching tasks. Our framework eliminates the need for complex hardware setups while enabling high-quality stereo image generation, making it valuable for both real-world applications and unsupervised learning scenarios. Project page is available at URL.
@@ 1 Introduction
Recent advancements in monocular depth estimation (MDE) models [32, 3, 61] have provided increasingly accurate dense disparity maps, playing a crucial role in image generation from single views. Meanwhile, Text-to-Image (T2I) diffusion models, such as Stable Diffusion [33], have made remarkable strides in generating diverse and high-quality images from user-provided text prompts. However, these models often struggle to maintain spatial consistency when altering the viewpoint of the generated images. While there has been some research on 3D novel view generation [51, 36], there remains a gap in utilizing image generation models to directly produce stereo images. StereoDiffusion [53] is the first approach to leverage diffusion-based generation for stereo image generation, but it struggles with pixel-level accuracy, as it applies disparity shifts in the latent space and fills occluded regions with blurry pixels that lack meaningful semantic content. On the other hand, Stable Diffusion Inpainting (SD-Inpainting) [33], though not designed for stereo image generation, faces similar challenges by incorrectly filling occluded areas with semantically inappropriate pixels.  Fig. 1 shows that previous diffusion-based methods and our methods tested on the COCO dataset [22]. The scale factor $\gamma$ is set to 0.15, meaning that the maximum disparity is 15% of the image width.
@@ 3 Methods
Figure 2: Overview of the GenStereo framework. ”+” denotes feature concatenation.
@@ 3.2 Cross-View Feature Enhancement
To facilitate effective information exchange between views, we adopt a two-parallel U-Nets framework pretrained from Stable Diffusion, inspired by GenWarp [36]. However, our work differs in its novel conditioning strategy: the reference U-Net takes concatenated $(I_{l},C_{l})$ as conditions to process the left view, while the denoising U-Net takes concatenated $(I_{\text{warp}},C_{r})$ as conditions when synthesizing the right view $\hat{I}_{r}$. This design ensures that each U-Net receives both image content and corresponding coordinate information for better feature extraction.
@@ 3.3 Training Strategy
To address dataset imbalance, we employ a resampling strategy. Specifically, we sample the smaller datasets multiple times until the number reaches $10\%$ of the largest dataset. As the Stable Diffusion model only accepts images of size $512\times 512$, we apply a random square crop and resize them to $512\times 512$.
@@ 4.1 Stereo Image Generation
Experimental Setup. We fine-tune the pretrained Stable Diffusion UNet v1.5 for 3 epochs across all experiments. For evaluation, we utilize two widely adopted stereo vision benchmarks: Middlebury 2014 [34] and KITTI 2015 [26], neither of which are included in our training set. Middlebury 2014 comprises 23 high-resolution indoor stereo pairs captured with wide-baseline stereo cameras under varying illumination conditions, providing challenging scenarios for stereo generation. KITTI 2015 contains 200 outdoor stereo pairs of street scenes with LiDAR-based sparse disparity ground truth, offering diverse real-world testing scenarios. Following StereoDiffusion [53], we preprocess Middlebury images to $512\times 512$ resolution, while KITTI images are center-cropped before resizing to $512\times 512$ to maintain the most informative regions.
@@
Quantitative Results. Tab. 2 presents the generation performance on Middlebury 2014 and KITTI 2015. The sparsity of ground truth disparities in KITTI limits evaluation accuracy, making pseudo-label disparities more effective for assessing stereo consistency. This sparsity particularly affects inpainting-based methods, as they struggle to infer accurate right-view images in areas with missing depth information, especially for dynamic objects and distant regions. In contrast, Middlebury 2014, which provides dense ground truth disparities, demonstrates a smaller performance gap between using ground truth and using pseudo disparities, underscoring the advantage of high-quality disparity annotations for generation. We obtain pseudo disparities from CREStereo [21] pretrained model, where evaluation datasets are not used for training to ensure unbiased evaluation.
GenStereo enables the construction of stereo datasets from a single camera, utilizing either sparse disparities or pseudo disparities. This capability expands the applicability of stereo generation to scenarios with limited sensor setups, making our approach more versatile in real-world applications where dual-camera systems may be impractical or cost-prohibitive.
@@
Methods Middlebury 2014 KITTI 2015
@@
Leave blank $11.328_{-3.489}^{+2.483}$ $0.315_{-0.154}^{+0.230}$ $0.450_{-0.089}^{+0.097}$ $12.980_{-3.831}^{+4.919}$ $0.374_{-0.251}^{+0.286}$ $0.313_{-0.109}^{+0.222}$
Stretch $14.842_{-2.753}^{+2.714}$ $0.432_{+0.265}^{-0.190}$ $0.285_{-0.089}^{+0.112}$ $14.757_{-4.694}^{+5.375}$ $0.429_{-0.271}^{+0.287}$ $0.212_{-0.100}^{+0.145}$
3D Photography [38] $14.190_{-2.798}^{+2.464}$ $0.427_{-0.175}^{+0.238}$ $0.275_{-0.065}^{+0.073}$ $14.540_{-4.023}^{+7.256}$ $0.398_{-0.323}^{+0.270}$ $0.210_{-0.073}^{+0.099}$
RePaint [24] $15.102_{-2.802}^{+2.909}$ $0.462_{-0.253}^{+0.268}$ $0.311_{-0.079}^{+0.088}$ $15.056_{-4.897}^{+5.366}$ $0.462_{-0.285}^{+0.268}$ $0.251_{-0.095}^{+0.128}$
@@
Table 3: Comparison with other diffusion-based stereo image generation methods for unsupervised stereo matching. The first group of methods is trained on PSMNet while the second group is trained on RAFT-Stereo.
@@
Table 4: Zero-Shot Generalization Benchmark. The first group shows methods trained with PSMNet, while the second group shows methods trained with RAFT-Stereo.
Methods KITTI 2012 KITTI 2015 ETH Middlebury-Q Middlebury-H
@@
NeRFStereo [48] 3.683 0.851 4.507 4.813 1.448 5.041 1.620 0.557 11.682 7.619 1.451 11.040 8.892 2.590 12.884
Ours 3.676 0.846 4.063 3.879 0.996 4.112 1.114 0.378 5.556 6.887 1.139 10.434 11.334 2.269 19.387
NeRFStereo [48] 3.787 0.855 4.140 5.356 1.469 5.561 0.830 0.301 3.370 5.819 1.003 8.370 7.426 1.820 10.948
Ours 3.224 0.734 3.455 4.683 1.042 4.897 0.959 0.294 2.667 3.450 1.003 7.952 4.446 1.514 9.938
@@ 4.2 Unsupervised Learning
We evaluate the performance of various unsupervised stereo matching approaches using the D1-all metric, which measures the percentage of incorrect disparity predictions (error $>3$px or $5\%$ of the true disparity), the End-Point-Error (EPE), and the >$\textgreater n$px metric, which denotes the percentage of pixels with an EPE exceeding $n$ pixels. To demonstrate the architecture-agnostic nature of our approach, we employ two distinct stereo models: PSMNet, a cost-volume-based method, and RAFT-Stereo, an iterative refinement model. For disparity estimation, we utilize the Depth Anything Model v2 (DAMv2) [61], from which we obtain the disparity map. The disparity values are normalized to the range $[0,1]$ and subsequently scaled by a factor $\gamma$ to enable flexibility generation.
@@
Table 5: Ablation studies.
Source # Mixed Datasets Random Drop Coord Embedding Warped Image Pixel Loss Fusion Middlebury 2014 KITTI 2015
@@ 5 Conclusion
Limitations. Despite promising results, diffusion-based models face inherent challenges with large disparities when generating right-view images due to large unconditioned regions. While our data augmentation such as random cropping and resizing mitigates these issues and performs well in typical stereo setups, future work could explore extending these models to accommodate larger disparities..

### v2

# Towards Open-World Generation of Stereo Images and Unsupervised Matching

@@ Abstract
Stereo images are fundamental to numerous applications, including extended reality (XR) devices, autonomous driving, and robotics. Unfortunately, acquiring high-quality stereo images remains challenging due to the precise calibration requirements of dual-camera setups and the complexity of obtaining accurate, dense disparity maps. Existing stereo image generation methods typically focus on either visual quality for viewing or geometric accuracy for matching, but not both. We introduce GenStereo, a diffusion-based approach, to bridge this gap. The method includes two primary innovations (1) conditioning the diffusion process on a disparity-aware coordinate embedding and a warped input image, allowing for more precise stereo alignment than previous methods, and (2) an adaptive fusion mechanism that intelligently combines the diffusion-generated image with a warped image, improving both realism and disparity consistency. Through extensive training on 11 diverse stereo datasets, GenStereo demonstrates strong generalization ability. GenStereo achieves state-of-the-art performance in both stereo image generation and unsupervised stereo matching tasks. Project page is available at https://qjizhi.github.io/genstereo.
@@ 1 Introduction
Recent advancements in monocular depth estimation (MDE) models [34, 3, 64, 65] have provided increasingly accurate dense disparity maps, playing a crucial role in image generation from single views. Meanwhile, Text-to-Image (T2I) diffusion models, such as Stable Diffusion [35], have made remarkable strides in generating diverse and high-quality images from user-provided text prompts. However, these models often struggle to maintain spatial consistency when altering the viewpoint of the generated images. While there has been some research on 3D novel view generation [53, 38], there remains a gap in utilizing image generation models to directly produce stereo images. StereoDiffusion [55] is the first approach to leverage diffusion-based generation for stereo image generation, but it struggles with pixel-level accuracy, as it applies disparity shifts in the latent space and fills occluded regions with blurry pixels that lack meaningful semantic content. On the other hand, Stable Diffusion Inpainting (SD-Inpainting) [35], though not designed for stereo image generation, faces similar challenges by incorrectly filling occluded areas with semantically inappropriate pixels.  Fig. 1 shows that previous diffusion-based methods and our methods tested on the COCO dataset [23]. The scale factor $\gamma$ is set to 0.15, meaning that the maximum disparity is 15% of the image width.
@@ 3 Methods
Figure 2: Overview of the GenStereo framework.
@@ 3.2 Cross-View Feature Enhancement
To facilitate effective information exchange between views, we adopt a two-parallel U-Nets framework pretrained from Stable Diffusion, inspired by GenWarp [38]. However, our work differs in its novel conditioning strategy: the reference U-Net takes concatenated $(I_{l},C_{l})$ as conditions to process the left view, while the denoising U-Net takes concatenated $(I_{\text{warp}},C_{r})$ as conditions when synthesizing the right view $\hat{I}_{r}$. This design ensures that each U-Net receives both image content and corresponding coordinate information for better feature extraction. As our model is conditioned on a reference image and its disparity map, we replace text condition of SD with image embedding of the left image through a CLIP image encoder in both U-Nets.
@@ 3.3 Training Strategy
To address dataset imbalance, we employ a resampling strategy. Specifically, we sample the smaller datasets multiple times until the number reaches $10\%$ of the largest dataset. As the Stable Diffusion model only accepts images of size $512\times 512$ for SD v1.5 and $768\times 768$ for SD v2.1, we apply a random square crop and resize.
@@ 4.1 Stereo Image Generation
Experimental Setup. We fine-tune the pretrained Stable Diffusion (SD) UNet for 3 epochs across all experiments. For evaluation, we utilize two widely adopted stereo vision benchmarks: Middlebury 2014 [36] and KITTI 2015 [28], neither of which are included in our training set. Middlebury 2014 comprises 23 high-resolution indoor stereo pairs captured with wide-baseline stereo cameras under varying illumination conditions, providing challenging scenarios for stereo generation. KITTI 2015 contains 200 outdoor stereo pairs of street scenes with LiDAR-based sparse disparity ground truth, offering diverse real-world testing scenarios. Following StereoDiffusion [55], we preprocess Middlebury images to $512\times 512$ resolution, while KITTI images are center-cropped before resizing to $512\times 512$ to maintain the most informative regions.
@@
Quantitative Results. Tab. 2 presents the generation performance on Middlebury 2014 and KITTI 2015. The sparsity of ground truth disparities in KITTI limits evaluation accuracy, making pseudo-label disparities more effective for assessing stereo consistency. This sparsity particularly affects inpainting-based methods, as they struggle to infer accurate right-view images in areas with missing depth information, especially for dynamic objects and distant regions. In contrast, Middlebury 2014, which provides dense ground truth disparities, demonstrates a smaller performance gap between using ground truth and using pseudo disparities, underscoring the advantage of high-quality disparity annotations for generation. We obtain pseudo disparities from CREStereo [22] pretrained model, where evaluation datasets are not used for training.
@@
Methods SD version Middlebury 2014 KITTI 2015
@@
Leave blank - $11.328_{-3.489}^{+2.483}$ $0.315_{-0.154}^{+0.230}$ $0.450_{-0.089}^{+0.097}$ $12.980_{-3.831}^{+4.919}$ $0.374_{-0.251}^{+0.286}$ $0.313_{-0.109}^{+0.222}$
Stretch - $14.842_{-2.753}^{+2.714}$ $0.432_{+0.265}^{-0.190}$ $0.285_{-0.089}^{+0.112}$ $14.757_{-4.694}^{+5.375}$ $0.429_{-0.271}^{+0.287}$ $0.212_{-0.100}^{+0.145}$
3D Photography [40] - $14.190_{-2.798}^{+2.464}$ $0.427_{-0.175}^{+0.238}$ $0.275_{-0.065}^{+0.073}$ $14.540_{-4.023}^{+7.256}$ $0.398_{-0.323}^{+0.270}$ $0.210_{-0.073}^{+0.099}$
RePaint [26] 1.5 $15.102_{-2.802}^{+2.909}$ $0.462_{-0.253}^{+0.268}$ $0.311_{-0.079}^{+0.088}$ $15.056_{-4.897}^{+5.366}$ $0.462_{-0.285}^{+0.268}$ $0.251_{-0.095}^{+0.128}$
SD-Inpainting [35] 1.5 $15.740_{-3.351}^{+3.511}$ $0.412_{-0.188}^{+0.236}$ $0.311_{-0.164}^{+0.098}$ $9.792_{-4.070}^{+2.809}$ $0.230_{-0.229}^{+0.236}$ $0.652_{-0.139}^{+0.144}$
StereoDiffusion [55] 1.5 $15.456_{-3.313}^{+2.669}$ $0.468_{-0.205}^{+0.252}$ $0.231_{-0.088}^{+0.096}$ $15.679_{-5.487}^{+5.888}$ $0.481_{-0.310}^{+0.245}$ $0.205_{-0.099}^{+0.135}$
@@
Table 3: Comparison with other diffusion-based stereo image generation methods for unsupervised stereo matching. The first group of methods is trained on PSMNet while the second group is trained on RAFT-Stereo. SD v1.5 is used for the experiments.
@@
Table 4: Zero-Shot Generalization Benchmark. The first group shows methods trained with PSMNet, the second group shows methods trained with RAFT-Stereo, and the third group represents IGEV++.
Methods KITTI 2012 KITTI 2015 ETH3D Middlebury-Q Middlebury-H
@@
NeRFStereo [50] 3.683 0.851 4.507 4.813 1.448 5.041 1.620 0.557 11.68 7.619 1.451 11.04 8.892 2.590 12.88
Ours + SD v1.5 3.676 0.846 4.063 3.879 0.996 4.112 1.114 0.378 5.556 6.887 1.139 10.43 11.33 2.269 19.39
Ours + SD v2.1 3.807 0.820 4.122 3.898 0.978 4.096 0.894 0.324 4.101 5.037 0.860 7.463 8.046 1.709 13.73
RAFT-Stereo [24] 3.944 0.830 4.239 5.254 1.132 5.468 0.883 0.267 2.607 5.157 1.217 10.25 5.774 1.436 11.20
NeRFStereo [50] 3.787 0.855 4.140 5.356 1.469 5.561 0.830 0.301 3.370 5.819 1.003 8.370 7.426 1.820 10.95
Ours + SD v1.5 3.224 0.734 3.455 4.683 1.042 4.897 0.959 0.294 2.667 3.450 1.003 7.952 4.446 1.514 9.94
Ours + SD v2.1 3.219 0.716 3.019 3.950 0.971 4.222 0.835 0.277 2.638 4.067 1.127 9.382 4.618 1.527 11.11
IGEV++ [62] 5.681 1.099 6.196 6.028 1.265 6.239 1.701 0.356 4.666 4.238 0.828 6.274 5.292 2.384 7.752
Ours + SD v2.1 2.802 0.686 3.015 3.746 0.952 4.004 1.226 0.326 3.429 3.682 0.684 5.512 2.797 0.842 5.098
@@ 4.2 Unsupervised Learning
We evaluate the performance of various unsupervised stereo matching approaches using the D1-all metric, which measures the percentage of incorrect disparity predictions (error $>3$px or $5\%$ of the true disparity), the End-Point-Error (EPE), and the >$\textgreater n$px metric, which denotes the percentage of pixels with an EPE exceeding $n$ pixels. To demonstrate the architecture-agnostic nature of our approach, we employ three distinct stereo models: PSMNet [6], a cost-volume-based method, RAFT-Stereo [24], an iterative refinement model, and IGEV++ [62], one of the most recent methods. For disparity estimation, we utilize the Depth Anything Model v2 (DAMv2) [65], from which we obtain the disparity map. The disparity values are normalized to the range $[0,1]$ and subsequently scaled by a factor $\gamma$ to enable flexibility generation.
@@
Table 5: Ablation studies. SD v1.5 is used for the experiments.
Disparity Source # Mixed Datasets Random Drop Coord Embedding Warped Image Pixel Loss Fusion Middlebury 2014 KITTI 2015
@@ 5 Conclusion
Limitations. Despite promising results, diffusion-based models face inherent challenges with large disparities when generating right-view images due to large unconditioned regions. While our data augmentation such as random cropping and resizing mitigates these issues and performs well in typical stereo setups, future work could explore extending these models to accommodate larger disparities.
Acknowledgments
We gratefully acknowledge the advanced computational resources provided by Engineering IT and Research Infrastructure Services at Washington University in St. Louis.